I need to solve for x in this equation: -3x+7=2y+3. What does x equal?
Answer by midwood_trail(310)

I need to solve for x in this equation: -3x+7=2y+3. What does x equal?
To solve for a variable means to isolate the variable.
In other words, we need to place x all alone on one side of the equation.
-3x + 7 = 2y + 3
Subtract 7 from both sides.
-3x = 2y + 3 - 7
-3x = 2y - 4
We now divide both sides by -3 to isolate x.
x = (2y - 4)/-3
Did you follow?